OpenStudy (anonymous):

Will someone tell me if i found the standard deviation correctly ? I will medal and fan . the data is 76 82 73 93 91 81 79 86 84

OpenStudy (anonymous):

First I found the mean which is 82.8 76+82+73+93+91+81+79+86+84=745 745/9 =82.8 Then I did (73-82.8)2 = (-9.8)2 = 96.04 (76-82.8)2 = (-6.8)2 = 46.24 (79-82.8)2 = (-3.8)2 = 14.44 (81-82.8)2 = (-1.8)2 = 3.24 (82-82.8)2 = (0.-8)2 = .64 (84-82.8)2 = (1.2)2 = 1.44 (86-82.8)2 = (3.2)2 = 10.24 (91-82.8)2 = (8.2)2 = 67.24 (93-82.8)2 = (10.2)2 = 104.04 I add all those up , and divide by (9-1) 96.04 + 46.24 + 14.44 + 3.24 + .64 + 1.44 + 10.24 + 67.24 + 104.04 = 343.56 343.56/8 = 42.95 , then take the square root of that . √42.95 = 6.55 the standard deviation is 6.55

OpenStudy (anonymous):

@AlexandervonHumboldt2 @Nnesha

Vocaloid (vocaloid):

yup, that's correct

OpenStudy (anonymous):

Really !

Vocaloid (vocaloid):

also, your work looks good to me

OpenStudy (anonymous):

thank you